### v1

Abstract
Virtual Try-On (trying clothes virtually) is a promising application of the Generative Adversarial Network (GAN). However, it is an arduous task to transfer the desired clothing item onto the corresponding regions of a human body because of varying body size, pose, and occlusions like hair and overlapped clothes. In this paper, we try to produce photo-realistic translated images through semantic segmentation and a generative adversarial architecture-based image translation network. We present a novel image-based Virtual Try-On application VTON-IT that takes an RGB image, segments desired body part, and overlays target cloth over the segmented body region. Most state-of-the-art GAN-based Virtual Try-On applications produce unaligned pixelated synthesis images on real-life test images. However, our approach generates high-resolution natural images with detailed textures on such variant images.  [^1] Details of the implementation, algorithms and codes, are publicly available on Github: https://github.com/shuntos/VITON-IT
keywords:
Virtual Try On, Human Part Segmentation, Image Translation, Semantic Segmentation, Generative Adversarial Network
1 Introduction
Research and Development on Virtual Try-On applications is getting popular as the fashion e-commerce market is rapidly growing. With a virtual try-on application, customers can try the desired cloth virtually before purchasing and sellers can benefit from an increased online marketplace. In addition, this application can reduce the uncertainty of size and appearance that most online shoppers are afraid of. Problems of traditional 3D-based virtual try-on are computational complexity, tedious hardware dependent data acquisition process and less user-friendly [1]. Image based 2D virtual try-on application, if integrated into existing e-commerce or digital marketplace, will be more scalable and memory sufficient as compared to 3D approach.
With the recent development in the GAN, Image-to-image translation on the conditional setting is possible [2]. Improved discriminator and generation architecture made cross-domain high-resolution image translation possible [3]. Which makes transformation of styles and textures from one domain to another possible.
In this paper, we mainly focus on solving the problems of 2D virtual try on applications by leveraging state-of-the-art deep learning network for semantic segmentation and robust image translation network for translating the input image into target domain. In previous works like VVT [4] there are problems with semantic segmentation because of plain backgrounds in training datasets. To solve this issue we trained a UNet-like semantic segmentation architecture on diverse images manually picked from the FGV6 dataset [5]. For image translation tasks residual mapping generator and the multi-scale discriminator is used which takes a semantic mask from the segmentation network and translates it into a wrapped RGB cloth with fine details. The previous methods only worked on an image with a single person on it [6]. Therefore, to make it applicable for multi-person cases, we used a pre-trained human detection Yolov5 model [7] trained on COCO dataset [8] to generate a bounding box for each human body and crop the overlaying cloth on it. VTON-IT architecture offers pose, background, and occlusion invariant application for the online fashion industry with a wide range of applications. Rigorous testing and experiments have shown that our approach generates more visually promising overlayed images as compared to existing methods.
2 Related Works
There are a number of image based virtual try-on approaches that have been tried in the prior researches, those relevant in our study are discussed here.
2.1 VITON
Han et al. presented an image-based Virtual Try-On Network (VITON) [9]: a coarse-to-fine framework that seamlessly transferred a target clothing item in a product image to the corresponding region of a clothed person in a 2D image. The warped target clothing-to match the pose of the clothed person was generated using a thin-plate spline (TPS) transformation which is ultimately fused with the person’s image.
2.2 CP-VTON
A similar structure to VITON is adopted by CPVTON [10], but it uses a neural network to learn the spatial transformation parameters of the TPS transformation present in its Geometric Matching Module (GMM). Therefore, the GMM will create a warped cloth image and a try-on module will fuse the warped cloth image to the target image of the person, preserving the precise features of the clothes.
2.3 CP-VTON+
CP-VTON+ [11] proposed a framework that preserves both cloth shape and texture through a two-stage architecture. The first stage is the Clothing Warping Stage, which transfers the texture of the clothing from the clothing image to the target person image. The later stage is the Blending Stage, which introduces a refinement module to further improve the quality of the generated try-on image. The framework consists of four major components: a body parsing network, a spatial transformer network, a shape transfer network, and a texture transfer network.
2.4 VTNFP
VTNFP [12] follows a three-stage design strategy by first generating warped clothing, followed by generating a body segmentation map of the person wearing the target clothing, and ending with a try-on synthesis module to fuse together all information for a final image synthesis. This method preserves target cloth and human body parts along with clothes that need no replacement.
2.5 Virtual Try-on auxiliary human segmentation
Virtual Try-On using auxiliary human segmentation [13] was based on additional improvements on existing CP-VTON. Using the human semantic segmentation prediction as an auxiliary task helped improve the virtual try-on performance. A branched architecture was proposed to simultaneously predict the try-on result and the expected segmentation mask of the generated try-on output where the target model was now wearing the in-shop cloth.
3 Proposed Approach
The network architecture proposed in this study is composed of three key components. The initial stage of the proposed architecture involves the implementation of a human body parsing network, which is followed by a body region segmentation network. Finally, an image translation network is employed to enable the wrapping of input clothing over the target body.
Figure 1: Proposed VITON-IT overview. First the human body is detected and cropped. Then, the desired body region is segmented through $U^{2}$-Net architecture and the segmented mask is fed to the image translation network to generate wrapped cloth. Finally, wrapped cloth is overlayed over the input image.
3.1 Human Parsing Network
To parse the body part for translation, the input image is given as input to Yolov5 pre-trained model trained on the Microsoft COCO dataset [8] for object detection tasks. This model outperforms existing object detection models in terms of latency and memory consumption. Model gives bounding boxes of humans in image and probability score. This allows the user to perform image translation on multiple objects. Pytorch implementation of the Yolov5-large model is used for inference. Input image of size $(640\times 640)$ is fed into the network with confidence threshold $=0.25$, Non-Maximum IOU threshold $=0.45$, and $max\_detection=10$ is used.
3.2 Human body Segmentation
$U^{2}$-Net architecture has been implemented to generate masks of desired body parts for image translation tasks. While existing backbones, such as Alexnet [14], VGG [15], ResNet [16], and DenseNet [17], are utilized for semantic segmentation tasks, their feature maps have a lower resolution. For example, ResNet reduces the size of feature maps to one-fourth of the input size. However, feature map resolution is very crucial for salient object detection (SOD). The objective of SOD is to segment the most visually appealing object in an image. Similarly, another problem of such backbones is their complex architecture because of additional feature extractor modules for extracting multi-level saliency features.
Nested UNet architecture, $U^{2}$-Net aims to go deeper while preserving feature map resolution. It can be trained from scratch and preserve the resolution of feature maps with the help of residual U-block (RSU). Top-level UNet-like architecture is followed by RSU which can extract Intra-stage multi-scale features.
Architecture: For salient object detection tasks, both local and global contextual information plays an important role. Existing feature extractor backbones have small receptive fields as they use small convolution filters of size $(1\times 1)$ or $(3\times 3)$ for the sake of computational cost and storage space. Such small receptive fields are not capable of capturing global information. Large receptive fields can capture global information from high-resolution feature maps. Dilated convolution operation results in an enlarged receptive field. However, performing multiple dilated convolution operations earlier with the original resolution needs high computational and memory resources. To tackle this problem RSU is used which has mainly three components.
  1. 1.
    Input Convolution layer: Transforms input feature $map(H,W,Cin)$ to intermediate feature map $F1(x)$ with $cout$. Local features are extracted from this plain convolution layer.
  2. 2.
    Input feature map $F1(x)$ is given as input to UNet like architecture of height $L$, higher value of $L$ means deeper the network with large number of pooling layers and ranges of receptive fields and numbers of local and global features. During downsampling multi-scale features are extracted and higher resolution features maps are encoded through progressive upsampling, concatenation and convolution.
  3. 3.
    Fusion of local and global features is done through RSU. $H(x)=(F(x)+U(F(x))$ Where $F(x)=$ Intermediate feature map, $U(F(x))=$ Multi-Scale contextual information and $H(x)=$ Desired mapping of input features.
$U^{2}$-Net consists of two-level nested U structures. The first level consists of 11 well-configured RSU stages capable of extracting intra-stage multi-scale and inter-stage multi-level features. Architecture has main 3 parts: a six-stage encoder, a five-stage decoder, and a saliency map fusion module with a decoder stage and the last encoder stage. In the first four encoder stages RSU is used but in the 5th and 6th stage resolution of the feature map is relatively low, so further downsampling might lead to loss of contextual information. Thus in the 5th and 6th encoder stages dilated RSU (replaced upsampling and pooling with dilated convolution) is used to preserve the resolution of feature maps. Feature map resolution in the 4th to 6th stages is the same. In decoder stages, dilated RSU is used when each stage takes concatenated upsampled feature maps from the previous stage. Saliency map fuser inputs saliency probability maps from five decoders and the last stage (6th) encoder where each map is generated through $(3\times 3)$ convolution and sigmoid function. These probability maps are concatenated and passed through $(1\times 1)$ convolution and sigmoid function to generate the final saliency probability map $S_{fused}$.
3.3 Image Translation
Masks generated by the human body segmentation network are passed through an image translation network (pix2pix). This network produces high-resolution photo-realistic synthesis RGB images from semantic label maps by leveraging a Generative Adversarial Network (GAN) in a conditional setting. Such visually appealing images are produced under adversarial training instead of any loss functions.
Architecture: This GAN framework consists of generator $G$ and discriminator $D$ for image-image translation tasks. The task of generator $G$ is to generate an image(RGB) of cloth given a binary semantic map generated by the human body segmentation network. Whereas, discriminator $D$ tries to classify whether the generated image is real or synthesized. Dataset consists of a pair of images ($S_{i}$ and $X_{i}$) where $S_{i}$ is a mask and $X_{i}$ is the corresponding real image. This architecture works in supervised configuration to model the conditional distribution of real images with given binary masks with a min-max game. Where generator $G$ and discriminator $D$ try to win against each other. From [18], we have,
$\min_{G}\max_{D}V(D,G)=\mathbb{E}_{\boldsymbol{x}\sim p_{\text{data }}(\boldsymbol{x})}[\log D(\boldsymbol{x})]+\mathbb{E}_{\boldsymbol{z}\sim p_{\boldsymbol{z}}(\boldsymbol{z})}[\log(1-D(G(\boldsymbol{z})))]$ (1)
This architecture uses UNet as a generator and a patch-based network as a discriminator. The generator takes 3 channel mask whereas a concatenated channel-wise semantic label map and the corresponding image are fed to Discriminator. The main component of this architecture is a coarse-to-fine generator, multi-scale discriminator, and optimized adversarial objective function. The generator consists of a global generator network $G1$ and a local enhancer network. Local generator outputs an image of resolution $4$ times larger than the original, or in other words ( 2 times of width and height) greater than the previous one. To increase the resolution of the synthesis image additional local generators can be added. For instance, the output resolution of $\{G1,G2)$ is $1024\times 2048$ whereas the output resolution of $\{G1,G2,G3\}$ is $2048\times 4096$. First global generator $G1$ is trained and then the local generator $G2$ is trained and so on, in the order of their resolution. Global generator $G1$ is trained on low-resolution images then another residual network $G2$ is added to $G1$ and a joint network is trained on higher-resolution images. Element wise sum of the feature map of $G2$ and the last feature map of $G1$ is fed into the next $G2$.
To differentiate real image and synthesis image, the discriminator must have a greater receptive field to capture global contextual information and similarly be able to extract lower-level local features. Discriminator architecture consists of $3$ different identical discriminators each working on different resolutions. The resolution of the real and synthesized image is downscaled by the factor of $2$ to create an image pyramid of $3$ scale. By this approach, a discriminator with the finest resolution helps the generator to produce an image with fine details.
$\min_{G}\max_{D_{1}D_{2}D_{3}}\sum_{k=1,2,3}\mathcal{L}_{GAN}\left(G,D_{k}\right)$ (2)
Where $D_{1}D_{2},D_{3}$ are $3$ scale discriminators [3]. As this architecture uses a multi-scale discriminator it extracts features from its multiple layers thus the generator has to produce a natural image at multiple scales. By adding such feature loss in the discriminator this helps to stabilize training loss.
Feature match loss is given by [3],
$\mathcal{L}_{GM}(G,D_{k})=\mathbb{E}_{(s,x)}\sum_{i=1}^{T}\frac{1}{N_{i}}[||D_{k}^{(i)}(\boldsymbol{s},\boldsymbol{x})-D_{k}^{(i)}(\boldsymbol{s},G(\boldsymbol{s}))||_{1}]$ (3)
Where $D_{k}$, $D_{k}^{(i)}$ are feature extractor layers of discriminator, $T$ represent total number of layers and $N$ represent number of elements per layer.
By adding feature match loss and GAN loss objective function becomes [3],
$min_{G}\left(\left(max_{D_{1},D_{2},D_{3}}\sum_{k=1,2,3}\mathcal{L}_{GAN}\left(G,D_{k}\right)\right)+\lambda\sum_{k=1,2,3}\mathcal{L}_{FM}(G,D_{k})\right)$ (4)
This loss function works well for translating mask to cloth image with higher resolution and detail texture.
Figure 2: Virtual Try-On Architecture: An input image is first fed to the YOLOv5 object detection model to detect the human body, which is then cropped. The cropped image is then passed through the $U^{2}$-Net segmentation model to generate a body region mask. Finally, the mask is fed into the Pix2Pix generator, which synthesizes RGB-based clothing onto the masked body region, resulting in a virtual try-on of the clothing. The output image shows the synthesized clothing on the original human body image.
4 Implementation Details
4.1 Training Human Body Segmentation network
For the training dataset, $6000$ good quality images are selected manually from FGVC6 [5] dataset and labeled using Labelme tool [19] to generate desired body masks. The average resolution of training images is $(630\times 1554)$. The model was trained through transfer learning using a pre-trained model trained on COCO dataset for general human body segmentation task with input Image size $320\times 320$ with random flip and crop. Pytorch library is used for training and inference. Adam optimizer [20] is used to train our network and its hyperparameters are set to default (initial learning rate $lr=1e-3$, $betas=(0.9,0.999)$, $eps=1e-8$, $weight\_decay=0$). And initially, the loss is set to $1$. The total number of iterations was $400000$ with a training loss of $0.109575$.
The model was trained on a custom dataset with unique ground truth. So performance was evaluated on a custom test dataset and we achieved $MaxFB=0.865$, Mean average Error $(MAE)=0.081$, $FBw=0.801$, and $SM=0.854$.
4.2 Training Image Translation Network
Dataset was prepared manually by creating a pair of real images and corresponding mask. Images are labeled using the Labelme tool to generate a semantic label and apply a series of geometrical augmentation algorithms. We prepared training pair images through data augmentation. As deep neural networks involve millions of parameters, incorporating more training data that is relevant to the domain can effectively mitigate the problem of overfitting. According to Zhao et. al [21], GAN performance is improved more by augmentations that cause spatial changes than by augmentations that just cause visual changes. Therefore, we have used several geometric augmentation techniques to augment the cloth image before training. These are namely Perspective Transform, Piecewise Affine Transform, Elastic Transformation, Shearing, and Scaling. We augmented both the image and its corresponding mask using the Imgaug library [22]. The image translation model was trained on a 3-channel input image of size $(512\times 512)$ with a batch size of $4$, without input label and instance map. The generator generates output images with the same number of image channels and shapes. Final loss consists of three components: GAN loss, discriminator-based feature matching loss, and VGG perceptual loss. After training up to $100$ epoch we got a GAN loss of $0.83$, a GAN Feature loss of $2.123$, and a VGG perceptual loss of $1.81$.
Figure 3: Example training image for human segmentation and image translation network
4.3 Training Setup
Both human segmentation and image translation network was trained on a Ubuntu 16.04 with 2 Nvidia GPU: GA102 [GeForce RTX 3090], 32 GB RAM and CPU: 20 core Intel(R) Core(TM) i9-10900K CPU @ 3.70GHz.
CP-VTON+ VTON-IT(proposed)
Reference Target Wrapped Final Wrapped Final
Image Clothes Clothes Result Clothes Result
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Visualized comparison with CP-VTON+
Figure 5: Result on outdoor images
5 Experimental Results
5.1 Qualitative Results
For evaluating the performance of VTON-IT through visual observation, we compared the final overlayed images with the output of CP-VTON+ [11]. Figure 4 shows that proposed virtual try-on application produces more realistic and convincing results in terms of texture transfer quality and pose preservation. Most of the existing virtual try on produce low-resolution output image. CP-VTON+ generates an output image with a fixed shape $(192\times 256)$ but our proposed approach works on high-resolution images. Through a high-resolution image translation network wrapped cloth of shape $(512\times 512)$ is generated. While experimenting with high-resolution input of shape $(2448\times 3264)$ we got a perfectly aligned natural-looking overlayed image with the same shape as input.
5.1.1 Result on Outdoor Images
Even though most of the images used for training body segmentation and translation are captured indoors with proper lighting conditions and predictable poses, both segmentation and translation models produce promising results on outdoor images with noisy backgrounds, unusual poses, and different lighting conditions. Figure 5 shows the results of the inference performed on outdoor images. However, the image on the right side has some artifacts in the unusual pose.
5.2 Quantitative Results
We adopted Structural Similarity Index(SSIM) [23], Multi-Scale Structural Similarity (MS-SSIM), Fréchet Inception Distance (FID) [24] and Kernel Inspection Distance (KID) [25] scores to measure the similarity between ground truths and synthesized images. The ground truths were made by manually wrapping clothes over the models’ images using various imaging tools and the synthesized images were the output of the model used. The result are shown in Table  1.
5.3 User Study
Although SSIM, MS-SSIM, FID, and KID can be used to determine the quality of image synthesis, it cannot reflect the overall realism and visual quality as assessed by human evaluation. Thus we performed a user study with 60 volunteers. In order to evaluate realism, volunteers were provided with two sets of images: ground truth images (manually wrapped clothes on the human models) and the outputs generated by our model. They were asked to score based on how real the clothes looked on the person and how well the texture of clothing was preserved. Then, they were asked to independently rate the photo-realism of only our output images. The result shows that our result was 70% similar to the ground truth and 60% photo-realistic.
Table 1: Quantitative evaluation of CP-VTON+ and VTON-IT in terms of SSIM, MS-SSIM, FID and KID scores.
Methods SSIM MS-SSIM FID KID
CPVTON+ 0.83 0.60 393 0.077
VTON-IT 0.93 0.87 50 0.019
6 Discussion
The effectiveness of this approach has been demonstrated through experiments conducted on both male and female bodies, under different lighting conditions, and with occlusions such as hands and hair, as well as varying poses. The first stage of the human detection network plays a crucial role in enhancing the overall performance of the application by reducing the input image size for later stages and eliminating unnecessary input regions by cropping the human body. To create an accurate segmentation map that adheres to geometric principles, it is necessary to train a human body segmentation network using ground truth images that contain detailed wrist and neck regions. Problem with public dataset like LVIS [26], MS COCO [8] and, Pascal Person Part dataset [27] is improper ground truth, hence we tried different pre-trained models like CDCL [28] , GRAPHONY [29], U2NET [30] to generate ground truth but none of these generate precise body mask. So, we manually picked 6000 good images of both men and women from the FGC6 dataset and labeled them manually. We trained a Generative image translation model on conditional settings with a pair of images as input. In order to generate a pair of images, we utilize geometrical augmentation techniques on both the semantic mask generated by the segmentation network and the corresponding real RGB image.
7 Conclusion and Future Works
The paper presents an innovative approach called VTON-IT (Virtual Try-On using Image Translation) that enables the transfer of desired clothing onto a person’s image, while accounting for variations in body size, pose, and lighting conditions. The proposed method outperforms existing approaches, as evidenced by quantitative and qualitative results that demonstrate the generation of natural-looking synthesized images. The proposed architecture consists of three components: human detection, body part segmentation, and image translation network. The image translation network is trained in this paper to generate synthesized images of the same domain (sweatshirts). However, it can be adapted for cross-domain synthesis by incorporating control parameters as label features. Future work may extend this approach to different types of clothing such as trousers, shorts, shoes, and more.
Acknowledgements: The authors would like to thank IKebana Solutions LLC for providing them with constant support for this research project.